# Comparing Task Graph Scheduling Algorithms: An Adversarial Approach

### v2

@@ 3. The SAGA Framework
in_trees randomly weighted in-trees randomly weighted
out_trees randomly weighted out-trees
chains randomly weighted parallel chains
blast synthetic Blast workflows (Hazekamp and Thain, 2017) Chameleon cloud inspired (Keahey et al., 2020)
bwa synthetic BWA workflows (Hazekamp and Thain, 2017)
cycles synthetic Cycles workflows (da Silva et al., 2019b)
epigenomics synthetic Epigenomics workflows (Juve et al., 2013)
genome synthetic 1000genome workflows (da Silva et al., 2019a)
montage synthetic Montage workflows (Rynge et al., 2014)
seismology synthetic Seismology workflows (Filgueira et al., 2016)
soykb synthetic SoyKB (Liu et al., 2016)
srasearch synthetic SRASearch workflows (Rynge, 2017)
etl RIoTBench ETL application (Shukla et al., 2017) Edge/Fog/Cloud Networks (Varshney et al., 2022)
predict RIoTBench PREDICT application (Shukla et al., 2017)
stats RIoTBench STATS application (Shukla et al., 2017)
train RIoTBench TRAIN application (Shukla et al., 2017)
@@
The in_trees, out_trees, and chains datasets each consist of $1000$ randomly generated network/task graph pairs following a common methodology used in the literature (Cordeiro et al., 2010). In-trees and out-trees are generated with between $2$ and $4$ levels (chosen uniformly at random), a branching factor of either $2$ or $3$ (chosen uniformly at random), and node/edge-weights drawn from a clipped gaussian distribution (mean: $1$, standard deviation: $1/3$, min: 00, max: $2$). Parallel chains task graphs are generated with between $2$ and $5$ parallel chains (chosen uniformly at random) of length between $2$ and $5$ (chosen uniformly at random) and node/edge-weights drawn from the same clipped gaussian distribution. Randomly weighted networks are complete graphs with between $3$ and $5$ nodes (chosen uniformly at random) and node/edge-weights drawn from the same clipped gaussian distribution.
@@
The scientific workflow datasets blast, bwa, cycles, epigenomics, genome, montage, seismology, soykb, and srasearch each contain $100$ problem instances. The task graphs are synthetically generated using the WfCommons Synthetic Workflow Generator (Coleman et al., 2023) and are based on real-world scientific workflows. The Chameleon cloud inspired networks are generated by fitting a distribution to the machine speed data from the execution traces (detailed information from a real execution of the application including task start/end times, cpu usages/requirements, data I/O sizes, etc.) of real workflows on Chameleon that are available in WfCommons and then sampling from that distribution to generate random networks. Because Chameleon uses a shared filesystem for data transfer, the communication cost can be absorbed into the computation cost and thus the communication strength between nodes is considered to be infinite.
@@
The IoT/edge-inspired etl, predict, stats, and train datasets each contain $1000$ problem instances. The task graphs and networks are generated using the approach described in (Varshney et al., 2022). The task graph structure is based on real-world IoT data streaming applications and the node weights are generated using a clipped gaussian distribution (mean: $35$, standard deviation: $25/3$, min: $10$, max: $60$). The input size of the application is generated using a clipped gaussian distribution (mean: $1000$, standard deviation: $500/3$, min: $500$, max: $1500$) and the edge weights are determined by the known input/output ratios of the tasks. The Edge/Fog/Cloud Networks are generated by constructing a complete graph with three types of nodes: edge nodes with CPU speed $1$, fog nodes with CPU speed $6$, and cloud nodes with CPU speed $50$. The communication strength between edge and fog nodes is $60$ and between fog and cloud/fog nodes is $100$. The number of edge, fog, and cloud nodes is between $75$ and $125$, $3$ and $7$, and $1$ and $10$, respectively (chosen uniformly at random). In order to generate a complete graph (a requirement for many of scheduling algorithms), the communication strength between edge and cloud nodes is set to $60$, the communication strength between cloud nodes is set to infinity (i.e., no communication delay).
@@
Many other schedulers and datasets have been proposed and used in the literature and can be easily integrated into SAGA in the future. SAGA is open-source (Authors, 2023a) and designed to be modular and extensible. We encourage the community to contribute new algorithms, datasets, and experimentation tools.
4. Benchmarking Results
@@
Figure 2. Makespan Ratios of 15 algorithms evaluated on 16 datasets. Gradients depict performance on different problem instances in each dataset.
@@
Figure 3. Comparison of Scheduling Algorithms on Slightly Modified Networks
@@
5. Adversarial Analysis
We propose a novel adversarial analysis method for comparing task scheduling algorithms called PISA (Problem-instance Identification using Simulated Annealing). The goal of this method is to identify problem instances on which one algorithm outperforms another. More formally, the goal is to find a problem instance $(N,G)$ that maximizes the makespan ratio of algorithm $\mathcal{A}$ against algorithm $\mathcal{B}$:
@@ 5. Adversarial Analysis
In doing so, we hope to fill in some of the gaps we know exist in the benchmarking results. We propose a simulated annealing-based approach for finding such problem instances. Simulated annealing is a meta-heuristic that is often used to find the global optimum of a function that has many local optima. In the context of our problem, the optimizer starts with an initial problem instance $(N,G)$ and then randomly perturbs the problem instance by changing the network or task graph in some way. If the perturbed problem instance has a higher makespan ratio than the current problem instance, the optimizer accepts the perturbation and continues the search from the new problem instance. If the perturbed problem instance has a lower makespan ratio than the current problem instance, the optimizer accepts the perturbation with some probability. This allows the optimizer to escape local optima and (potentially) find the global optimum. Over time, the probability of accepting perturbations that decrease the makespan ratio decreases, allowing the optimizer to settle into a high-makespan ratio state. The pseudocode for our proposed simulated annealing process is shown in Algorithm 1.
Algorithm 1 PISA (Problem-instance Identification using Simulated Annealing)
1:procedure PISA($N,G,T_{max},T_{min},I_{max},\alpha,\mathcal{A},\mathcal{B}$)
2:$(N,G)$ is the initial problem instance
3:$T_{max}$ is the initial (maximum) temperature
4:$T_{min}$ is the minimum temperature
5:$I_{max}$ is the maximum number of iterations
6:$\alpha$ is the cooling rate
7:$\mathcal{A}$ is the scheduler
$\mathcal{B}$ is the baseline scheduler
9:     $N_{best}\leftarrow N$
10:     $G_{best}\leftarrow G$
11:     $M_{best}\leftarrow\frac{M_{\mathcal{A}(N,G)}}{M_{\mathcal{B}(N,G)}}$
12:     $T\leftarrow T_{max}$
13:     $i\leftarrow 0$
14:     while $T>T_{min}$ and $i<I_{max}$ do
15:         $N^{\prime},G^{\prime}\leftarrow$ Perturb($N,G$)
16:         $M^{\prime}\leftarrow\frac{M_{\mathcal{A}(N^{\prime},G^{\prime})}}{M_{\mathcal{B}(N^{\prime},G^{\prime})}}$
17:         if $M^{\prime}>M_{best}$ then
18:              $N_{best}\leftarrow N^{\prime}$
19:              $G_{best}\leftarrow G^{\prime}$
20:              $M_{best}\leftarrow M^{\prime}$
21:         else
22:              $p\leftarrow\exp\left(-\frac{M^{\prime}/M_{best}}{T}\right)$
23:              $r\leftarrow$ Random($0,1$) $\triangleright$ Uniform random number between 00 and $1$
24:              if $r<p$ then
25:                  $N_{best}\leftarrow N^{\prime}$
26:                  $G_{best}\leftarrow G^{\prime}$
27:                  $M_{best}\leftarrow M^{\prime}$
28:              end if
29:         end if
30:         $T\leftarrow T\cdot\alpha$
31:         $i\leftarrow i+1$
32:     end while
33:     return $N_{best},G_{best},M_{best}$
34:end procedure
The Perturb function is responsible for perturbing the problem instance by changing the network or task graph in some way. In our implementation, the Perturb function randomly selects (with equal probability) one of the following perturbations:
  1. (1)
@@
  2. (2)
    Change Network Edge Weight: Same as Change Network Node Weight, but for edges (not including self-edges).
  3. (3)
@@ 5.1. Results
Figure 4. Heatmap of Makespan Ratios for all 15 Algorithms compared to each other. The cell value (and color) for row $i$ and column $j$ indicates the Makespan Ratio for the worst-case instance found by PISA for scheduler $j$ against baseline scheduler $i$.
@@
The benefit of our approach is clear from just a quick glance at Figure 4 (there’s a lot more red!). A closer look reveals even more interesting results. First, PISA finds, for every scheduling algorithm, problem instances such that it performs at least twice as bad as another of the algorithms. In fact, for most of the algorithms ($10$ of 15), it finds a problem instance such that the algorithm performs at least five times worse than another algorithm! Furthermore, we note that for nearly every pair of algorithms $\mathcal{A},\mathcal{B}$, PISA identifies a problem instance where $\mathcal{A}$ outperforms $\mathcal{B}$ and where $\mathcal{B}$ outperforms $\mathcal{A}$. Finally, some of the cells in Figure 4 have values of $>1000$. In this case, PISA identified a problem instance where one algorithm drastically outperforms the other. For these cases, it’s likely that there exist problem instances where the scheduler performs arbitrarily poorly compared to the baseline scheduler.
Readers can find a closer look at one example comparing HEFT and CPoP (two well-known task scheduling algorithms (Topcuoglu et al., 1999)) in Appendix A.2 that demonstrates how PISA can be used to discover the reasons why an algorithm performs well on some problem instances and poorly on others.
6. Application-Specific PISA
@@
In Section 5, we introduced PISA as an effective method for finding problem instances where an algorithm performs much worse than benchmarking results suggest. The results, however, depend greatly on the initial problem instance and the implementation of the Perturb function. These two things define the space of problem instances the algorithm searches over and also affect which problem instances are more or less likely to be explored. We chose an implementation in Section 5 that kept problem instances relatively small (between three and five tasks and compute nodes) and allowed arbitrary task graph structure and CCRs (communication-to-computation ratios). By keeping the problem instance size small but allowing for almost arbitrary task graph structure and CCR, this implementation allowed us to explore how the structure of problem instances affects schedules in order to find patterns where certain algorithms out-perform others. In many more realistic scenarios, though, application developers have a better idea of what their task graph and/or compute network will look like. PISA can be easily restricted to searching over a space of realistic problem instances by adjusting the Perturb implementation and initial problem instance. In this section, we report results on experiments with application-specific Perturb implementations. Again, these results support the main hypothesis of this paper that PISA reveals performance boundaries that a traditional benchmarking approach does not.
6.1. Experimental Setup

### v3

@@ IV-B Datasets
in_trees in-trees randomly weighted
out_trees out-trees
chains parallel chains
blast Blast workflows [25] Chameleon cloud inspired [26]
bwa BWA workflows [25]
cycles Cycles workflows [27]
epigenomics Epigenomics workflows [28]
genome 1000genome workflows [29]
montage Montage workflows [30]
seismology Seismology workflows [31]
soykb SoyKB [32]
srasearch SRASearch workflows [33]
etl IoT ETL application [34] Edge, Fog, & Cloud Networks [35]
predict IoT PREDICT application [34]
stats IoT STATS application [34]
train IoT TRAIN application [34]
@@
The in_trees, out_trees, and chains datasets each consist of $1000$ randomly generated network/task graph pairs following a common methodology used in the literature [36]. In-trees and out-trees are generated with between $2$ and $4$ levels (chosen uniformly at random), a branching factor of either $2$ or $3$ (chosen uniformly at random), and node/edge-weights drawn from a clipped gaussian distribution (mean: $1$, standard deviation: $1/3$, min: 00, max: $2$). Parallel chains task graphs are generated with between $2$ and $5$ parallel chains (chosen uniformly at random) of length between $2$ and $5$ (chosen uniformly at random) and node/edge-weights drawn from the same clipped gaussian distribution. Randomly weighted networks are complete graphs with between $3$ and $5$ nodes (chosen uniformly at random) and node/edge-weights drawn from the same clipped gaussian distribution.
@@
The scientific workflow datasets blast, bwa, cycles, epigenomics, genome, montage, seismology, soykb, and srasearch each contain $100$ problem instances. The task graphs are synthetically generated using the WfCommons Synthetic Workflow Generator [37] and are based on real-world scientific workflows. The Chameleon cloud inspired networks are generated by fitting a distribution to the machine speed data from the execution traces (detailed information from a real execution of the application including task start/end times, cpu usages/requirements, data I/O sizes, etc.) of real workflows on Chameleon that are available in WfCommons and then sampling from that distribution to generate random networks. Because Chameleon uses a shared filesystem for data transfer, the communication cost can be absorbed into the computation cost and thus the communication strength between nodes is considered to be infinite.
@@
The IoT/edge-inspired etl, predict, stats, and train datasets each contain $1000$ problem instances. The task graphs and networks are generated using the approach described in [35]. The task graph structure is based on real-world IoT data streaming applications and the node weights are generated using a clipped gaussian distribution (mean: $35$, standard deviation: $25/3$, min: $10$, max: $60$). The input size of the application is generated using a clipped gaussian distribution (mean: $1000$, standard deviation: $500/3$, min: $500$, max: $1500$) and the edge weights are determined by the known input/output ratios of the tasks. The Edge/Fog/Cloud Networks are generated by constructing a complete graph with three types of nodes: edge nodes with CPU speed $1$, fog nodes with CPU speed $6$, and cloud nodes with CPU speed $50$. The communication strength between edge and fog nodes is $60$ and between fog and cloud/fog nodes is $100$. The number of edge, fog, and cloud nodes is between $75$ and $125$, $3$ and $7$, and $1$ and $10$, respectively (chosen uniformly at random). In order to generate a complete graph (a requirement for many scheduling algorithms), edge-cloud and cloud-cloud communication strengths are set to $60$ and infinity (i.e., no communication delay), respectively.
@@
Many other schedulers and datasets have been proposed and used in the literature and can be easily integrated into SAGA in the future. SAGA is open-source [3] and designed to be modular and extensible. We encourage the community to contribute new algorithms, datasets, and tools.
V Benchmarking Results
@@
Figure 2: Makespan Ratios of 15 algorithms evaluated on 16 datasets. Gradients depict performance on different problem instances in each dataset.
@@
Figure 3: Comparison of Scheduling Algorithms on Slightly Modified Networks
@@
VI Adversarial Analysis
We propose a novel adversarial analysis method for comparing task scheduling algorithms called PISA (Problem-instance Identification using Simulated Annealing). The goal of this method is to identify problem instances on which one algorithm maximally outperforms another. More formally, the goal is to find a problem instance $(N,G)$ that maximizes the makespan ratio of algorithm $\mathcal{A}$ against algorithm $\mathcal{B}$:
@@ VI Adversarial Analysis
In doing so, we hope to fill in some of the gaps we know exist in the benchmarking results. We propose a simulated annealing-based approach for finding such problem instances. Simulated annealing is a meta-heuristic that is often used to find the global optimum of a function that has many local optima. In the context of our problem, the optimizer starts with an initial problem instance $(N,G)$ and then randomly perturbs the problem instance by changing the network or task graph in some way. If the perturbed problem instance has a higher makespan ratio than the current problem instance, the optimizer accepts the perturbation and continues the search from the new problem instance. If the perturbed problem instance has a lower makespan ratio than the current problem instance, the optimizer still accepts the perturbation with some probability. This allows the optimizer to escape local optima and (potentially) find a global optimum. Over time, the probability of accepting perturbations that decrease the makespan ratio decreases, allowing the optimizer to settle into a high-makespan ratio state before terminating. The pseudocode for our proposed simulated annealing process is presented in Algorithm 1.
Algorithm 1 PISA
1:Initialize solution $(N,G)$ and best solution $(N_{best},G_{best})$
2:Set initial temperature $T=T_{max}$
3:while $T>T_{min}$ and iteration $<I_{max}$ do
4:     Generate new candidate $(N^{\prime},G^{\prime})$ by perturbing $(N,G)$
5:     Evaluate new solution’s quality $M^{\prime}$
6:     if $M^{\prime}>M_{best}$ then
7:         Update best solution $(N_{best},G_{best})$
8:     else
9:         Accept with probability $\exp\left(-\frac{M^{\prime}/M_{best}}{T}\right)$
10:     end if
11:     Reduce temperature: $T=T\cdot\alpha$
12:end while
13:return best solution $(N_{best},G_{best})$
In our implementation, we perturb problem instances by randomly selecting (with equal probability) one of the following perturbations:
  1. 1.
@@
  2. 2.
    Change Network Edge Weight: Same as Change Network Node Weight, but for (non-self) edges.
  3. 3.
@@ VI-A Results
Figure 4: Heatmap of makespan ratios for all 15 Algorithms compared to each other. The cell value (and color) for row $i$ and column $j$ indicates the makespan ratio for the worst-case instance found by PISA for scheduler $j$ against scheduler $i$.
@@
The benefit of our approach is clear from just a quick glance at Figure 4 (there’s a lot more red!). A closer look reveals even more interesting results. First, PISA finds, for every scheduling algorithm, problem instances on which it performs at least twice as bad as another of the algorithms. In fact, for most of the algorithms ($10$ of 15), it finds a problem instance such that the algorithm performs at least five times worse than another algorithm. For HEFT, one of the most popular scheduling algorithms, PISA finds a problem instance where it performs $4.34$ times worse than FastestNode, a (generally poor-performing) simple baseline algorithm that schedules all tasks to execute in serial on the fastest compute node. In fact, for many of the algorithms, PISA finds problem instances where they perform quite poorly compared to FastestNode, perhaps suggesting that a common weakness of many scheduling algorithms is over-parallelization on some problem instances.
Furthermore, we note that for nearly every pair of algorithms $\mathcal{A},\mathcal{B}$, PISA identifies a problem instance where $\mathcal{A}$ outperforms $\mathcal{B}$ and where $\mathcal{B}$ outperforms $\mathcal{A}$. In other words, we don’t see many algorithms that are strictly better or worse than others. Each algorithm has its own strengths and weaknesses that are revealed by PISA. Finally, some of the cells in Figure 4 have values of $>1000$. In this case, PISA identified a problem instance where one algorithm drastically outperforms the other. For these cases, it’s likely that there exist problem instances where the scheduler performs arbitrarily poorly compared to the baseline scheduler. Insights like these (of which Figure 4 contains many more), provided by PISA, are invaluable for understanding the strengths and weaknesses of task scheduling algorithms and can be used to guide the development of new algorithms and distributed computing systems.
VI-B Case Study: HEFT vs. CPoP
The HEFT and CPoP algorithms were proposed in the same paper by Topcuoglu et al. in 1999 and have remained two of the most popular task scheduling algorithms in decades since. Both are list-scheduling algorithms that use a priority list to determine the order in which tasks are greedily scheduled. In HEFT, each task is scheduled on the node that minimizes its finish time, given previous scheduling decisions. CPoP is similar, but commits to scheduling all critical-path tasks (those on the longest path in the task graph) on the fastest node [^3] This statement is true for the related machines model. In general, CPoP schedules critical-path tasks to the node that minimizes the sum of execution times of critical-path tasks on that node.. The priority functions for these algorithms are slightly different from each other, but are both based on task distance (in terms of average compute and communication time) from the start and/or end of the task graph. In HEFT, the priority of a sink task (one with no dependencies) is its average execution time over all nodes in the network. A non-sink task’s priority is the sum of its average execution time over all nodes in the network plus the maximum communication and compute time of its successors. In other words, it’s the average amount of time it takes to execute the task on any node in the network plus the maximum amount of time it takes (in an ideal scenario) to execute the rest of the task graph. In CPoP, the task priorities are computed based on the distance to the end and from the start of the task graph. This difference is a large part of what causes the two algorithms to perform differently on certain problem instances. Observe in Figure 5 a problem instance identified by PISA where HEFT performs approximately $1.55$ times worse than CPoP.
(a) Task Graph
(b) HEFT Schedule
(c) Network
(d) CPoP Schedule
Figure 5: Problem Instance where HEFT performs $\approx 1.55$ times worse than CPoP.
In this instance, the algorithms differ primarily in whether task $A$ or task $C$ has higher priority. In both algorithms, task $B$ has the highest priority and is thus scheduled first on node $2$ (the fastest node). For CPoP, task $C$ must have the highest priority because it is on the critical path $B\rightarrow C$. For HEFT, though, task $A$ has the higher priority than $C$ because it is further from the end of the task graph. As a result, CPoP schedules task $C$ on node $2$, which allows task $A$ to be scheduled and executed in parallel on node $3$ (the second fastest node). HEFT, on the other hand, schedules all tasks on node $2$ and thus does not benefit from parallel execution. CPoP succeeds in this instance because it prioritizes tasks that are on the critical path, keeping high-cost tasks/communications on the same node and allowing low-cost tasks/communications to be executed in parallel on other nodes. HEFT greedily schedules high-cost tasks on fast nodes without taking into as much consideration how doing so might affect the rest of the task graph (especially the tasks on the critical path).
This does not mean that CPoP is better than HEFT, though. Observe in Figure 6 a problem instance identified by PISA where CPoP performs approximately $2.83$ times worse than HEFT.
(a) Task Graph
(b) HEFT Schedule
(c) Network
(d) CPoP Schedule
Figure 6: Problem Instance where CPoP performs $\approx 2.83$ times worse than HEFT.
The problem that CPoP faces for this instance is exactly what allows it to succeed in the previous instance: committing to scheduling all critical-path tasks on the fastest node. In this instance, the critical path is $B\rightarrow C$. Thus, CPoP schedules task $C$ on node $2$ (the fastest node) even though it would have finished much faster (due to the communication cost incurred by its dependency on task $A$) on node $1$. HEFT does not have this problem because it does not commit to scheduling all critical-path tasks on the fastest node.
One of the most promising aspects of PISA is that it allows us to identify patterns where certain algorithms outperform others. We can generalize the patterns we observe in the above examples to formalize a set of conditions under which CPoP outperforms HEFT and vice versa. For example, we might hypothesize that HEFT will perform poorly against CPoP for fork-join task graphs where one chain of tasks has a much higher initial communication cost than the other. Formally, let us define a set of task graphs with tasks $A,B,C,D$ such that tasks $A$ and $D$ have cost $1$, tasks $B$ and $C$ have random cost (drawn from a clipped gaussian distribution with mean $10$, standard deviation $10/3$, and min 00), and the dependencies $A\rightarrow B$, $A\rightarrow C$, and $B\rightarrow D$ have cost $1$ and $C\rightarrow D$ has high cost (drawn from a clipped gaussian distribution with mean $100$, standard deviation $100/3$, and min 00). Observe in Figure 7 an example of a task graph from this set and the makespans of both HEFT and CPoP on $1000$ randomly generated instances from this set (on a completely homogeneous network, for simplicity).
(a) Task Graph
(b) Makespans
Figure 7: Task Graph and Makespans for HEFT and CPoP on a set of task graphs where HEFT performs poorly.
Observe that HEFT performs significantly worse than CPoP on this family of task graphs.
On the other hand, we might hypothesize that CPoP will perform poorly against HEFT for task graphs where initially low communication costs guide CPoP into parallelizing tasks poorly and force it into a situation where it must perform a high-cost communication. Formally, let us define a set of task graphs with tasks $A$ through $K$ such that task $A$ is the start task, tasks $B$ through $J$ depend on task $A$, and task $K$ depends on tasks $B$ through $J$. All tasks have random costs drawn from a clipped gaussian distribution with mean $1$ and standard deviation $1/3$. The dependencies from task $A$ to the inner tasks $B$ through $J$ have random costs drawn from a clipped gaussian distribution with mean $1$ and standard deviation $1/3$ while the dependencies from the inner tasks to task $K$ have random costs drawn from a clipped gaussian distribution with mean $10$ and standard deviation $10/3$. In other words, the communication on the “join” part of the task graph is ten times more expensive than the communication on the “fork” part of the task graph. Let the network be a complete graph with four nodes where the fastest node has a weak link to the second fastest node. Let the fastest node have a speed of $3$ and the speeds of the other nodes be drawn from a clipped gaussian distribution with mean $1$ and standard deviation $1/3$ (they are, on average, three times slower than the fastest node). Let the communication costs between the fastest and second fastest node be drawn from a clipped gaussian distribution with mean $1$ and standard deviation $1/3$ and the communication costs between all other nodes be drawn from a clipped gaussian distribution with mean $10$ and standard deviation $5/3$ (the communication link between the two fastest nodes is on average ten times weaker than the communication links between all other nodes). Observe in Figure 8 an example of a task graph from this set and the makespans of both HEFT and CPoP on $1000$ randomly generated instances from this set.
(a) Task Graph
(b) Makespans
Figure 8: Task Graph and Makespans for HEFT and CPoP on a set of task graphs where CPoP performs poorly.
CPoP performs significantly worse than HEFT on this family of task graphs.
By identifying problem instances where an algorithm performs significantly worse than another, PISA allows us to gain insight into the strengths and weaknesses of different algorithms. Performing this kind of analysis on other pairs of algorithms remains an important task for future work.
VII Application-Specific PISA
@@
In Section VI, we introduced PISA as an effective method for finding problem instances where an algorithm performs much worse than benchmarking results suggest. The results, however, depend greatly on the initial problem instance and the implementation of the Perturb function. These two things define the space of problem instances the algorithm searches over and also affect which problem instances are more or less likely to be explored. We chose an implementation in Section VI that kept problem instances relatively small (between three and five tasks and compute nodes) and allowed arbitrary task graph structure and CCRs (communication-to-computation ratios). By keeping the problem instance size small but allowing for almost arbitrary task graph structure and CCR, this implementation allowed us to explore how the structure of problem instances affects schedules in order to find patterns where certain algorithms out-perform others. In many more realistic scenarios, though, application developers have a better idea of what their task graph and/or compute network will look like. PISA can be easily restricted to searching over a space of realistic problem instances by adjusting the Perturb implementation and initial problem instance. In this section, we report results on experiments with application-specific Perturb implementations. Again, these results support the main hypothesis of this paper that PISA reveals performance boundaries that a traditional benchmarking approach does not.
VII-A Experimental Setup